# Solving Fredholm Integral Equations of the Second Kind via Wasserstein Gradient Flows

### v2

@@ 2.1 Stability of $\mathcal{F}_{\alpha}$
Assumption 1 restricts the class of kernels $\mathrm{k}$ to smooth kernels with bounded derivatives up to the third order. This might seem restrictive, however, smooth kernels are particularly challenging while discountinuous or degenerate kernels normally allow for specific solution methodologies (see, e.g., [40, Ch. 11]). In this section, we will require only that $\mathrm{k}\in\mathrm{C}(\mathbb{R}^{d}\times\mathbb{R}^{d},[0,+\infty))$. The control of higher order derivatives is used to establish convergence result for the interacting particle system (in both continuous and discrete time) .
@@ 2.2 Convergence of Minimizers
We now consider the limiting behaviour of the regularized functional $\mathcal{F}_{\alpha}^{\eta}$ when the parameters $\alpha,\eta$ converge to 0. In particular, we show that if both parameters tend to 0 the minimizer of $\mathcal{F}_{\alpha}^{\eta}$ converges to one of the minimizers of the unregularized functional.
@@
Under this assumption we have the following results which describe the behaviour of the minimizer as $\eta\to 0$ and as $(\alpha,\eta)\to(0,0)$ jointly.
@@
Assume 1 and 2.
@@
    If there exist $\pi_{\alpha}^{\star}\in\mathcal{P}_{2}^{ac}(\mathbb{R}^{d})$, $(\eta_{n})_{n\in\mathbb{N}^{\star}}\in[0,+\infty)^{\mathbb{N}^{\star}}$ such that $\lim_{n\to+\infty}\eta_{n}=0$ and $\lim_{n\to+\infty}\mathbf{W}_{2}(\pi_{\alpha,\eta_{n}}^{\star},\pi_{\alpha}^{\star})=0$, then
    $\pi_{\alpha}^{\star}\in\arg\min_{\mathcal{P}_{2}^{ac}(\mathbb{R}^{d})}\mathcal{F}_{\alpha}(\pi).$
@@
    If there exist $\pi^{\star}\in\mathcal{P}_{2}^{ac}(\mathbb{R}^{d})$, $(\eta_{n})_{n\in\mathbb{N}^{\star}}\in[0,+\infty)^{\mathbb{N}^{\star}},(\alpha_{n})_{n\in\mathbb{N}^{\star}}\in(0,+\infty)^{\mathbb{N}}$ such that $\lim_{n\to+\infty}\eta_{n}=0,\lim_{n\to+\infty}\alpha_{n}=0$ and $\lim_{n\to+\infty}\mathbf{W}_{2}(\pi_{\alpha,\eta_{n}}^{\star},\pi^{\star})=0$, then
    $\pi^{\star}\in\arg\min_{\mathcal{P}_{2}^{ac}(\mathbb{R}^{d})}\mathcal{F}(\pi).$
@@ 2.3 Maximum Entropy Methods
Maximum entropy solutions to Fredholm integral equations were originally proposed in [45] under moment constraints obtained by selecting different families of basis functions [36, 33]. In the one dimensional case, such a solution can be written analytically (see, e.g., [33, Prop. 3.1]), and these approaches are particularly effective. However, in higher dimensions maximum entropy methods generally require discretization of the support of the solution $\pi$.
@@ 3.2 McKean–Vlasov SDE
The previous result deals with the behaviour of $\pi_{t}$ for fixed $\alpha,\eta$. Under our assumptions, [14, Thm. 4.1] and [47, Thm. 1] guarantee that $\mathcal{F}^{\eta}(\pi_{t})-\inf\mathcal{F}^{\eta}\to 0$ provided that the parameter $\alpha$ decays at rate $1/\log t$.
@@ 5.1 Simple Gaussian Example
We start by considering a simple analytically-tractable integral equation. Consider, for $\lambda\in(0,1)$,  (1) with $\varphi(x)=(1-\lambda)\mathcal{N}\left(x;0,1\right)$ and $\mathrm{k}(x,y)=\mathcal{N}\left(y;xe^{-\beta},1-e^{-2\beta}\right)$ for some $\beta>0$, with $\mathcal{N}(x;\mu,\sigma^{2})$ denoting the density of a normal distribution of mean $\mu$ and variance $\sigma^{2}$ evaluated at $x$. Since $\max_{x}\int_{\mathbb{R}^{d}}\lambda k(x,y)\mathrm{d}y=\lambda<1$, [40, Cor. 2.16] guarantees that (1) admits a unique solution, $\pi(x)=\mathcal{N}\left(x;0,1\right)$ for all $0<\lambda<1$. In our experiments we set $\beta=0.5$.
@@ 5.1.2 Comparison with Reversible jump MCMC
The previous experiments show that the presence of the reference measure significantly improves the results provided by FE2kind-WGF. We now compare our approach with the MCMC algorithm of [23] which provides a Monte Carlo approximation of the von Neumann series representation of the solution $\pi$ (see, e.g., [40, Sec. 2.4]). We implement their reversible jump MCMC (RJ-MCMC) with the following set up: we set the death and birth probabilities to 1/3 and use a random walk proposal with variance $\sigma^{2}=0.1^{2}$ for the update move and sample from $\pi$ for the birth move.
@@ 5.2 Karhunen–Loève Expansions
The Karhunen–Loève expansion of a given stochastic process with covariance function $\mathrm{k}(x,y)$ is given by the pairs of eigenvalues and eigenfunctions $(\lambda,\pi)$ of $\mathrm{k}$. For any eigenvalue $\lambda$, the corresponding eigenfunction can be found by solving (1) with $\varphi\equiv 0$. We consider two examples commonly used in the Gaussian process literature [60, Sec. 4.2]: the squared exponential kernel, $\mathrm{k}(x,y)=\exp(-(y-x)^{2})$, which satisfies our assumptions, and the exponential kernel, $\mathrm{k}(x,y)=\exp(-|y-x|)$, which is not differentiable for $y=x$ and therefore does not satisfy Assumption 1.
@@ A.2 Proof of Proposition 1
    $\mathcal{F}^{\eta}(\pi)\geq$ $-e^{-1}+\frac{\eta}{(1+\lambda)\mathtt{M}+\eta}\mathcal{F}^{\eta}(\pi)\geq-\left(1-\frac{\eta}{(1+\lambda)\mathtt{M}+\eta}\right)^{-1}e^{-1}=:\mathtt{C}$
@@
    $\mathcal{F}_{\alpha}^{\eta}(\pi_{0})=-\mathrm{H}(\pi_{0})-\int_{\mathbb{R}^{d}}\pi_{0}(x)\log(\varphi(x)+\pi_{0}[\mathrm{k}(x,\cdot)]+\eta)\mathrm{d}x.$
@@
    $|\log(\varphi(x)+\pi[\mathrm{k}(x,\cdot)]+\eta)|\leq\max(|\log(2\mathtt{M}+\eta)|,|\log(\eta)|),$
@@
    $-\int_{\mathbb{R}^{d}}\pi_{0}(x)\log(\varphi(x)+\pi_{0}[\mathrm{k}(x,\cdot)]+\eta)\mathrm{d}x\leq\max(|\log(2\mathtt{M}+\eta)|,|\log(\eta)|).$
@@ B.2 Proof of Proposition 2
Since $t\mapsto t\log(t)$ is bounded on $[0,1]$ we have that
@@
    By Lemma 2, $\pi_{\alpha,\eta_{n}}^{\star}\in\mathcal{P}_{2}^{ac}(\mathbb{R}^{d})$. Consider for fixed $\alpha$ the family $\{\mathcal{F}_{\alpha}^{\eta}:\eta\geq 0\}$. Let $(\eta_{n})_{n\in\mathbb{N}^{\star}}\in[0,+\infty)^{\mathbb{N}^{\star}}$ be such that $\underset{n\to+\infty}{\lim}\eta_{n}=0$ and $\underset{n\to+\infty}{\lim}\pi_{\alpha,\eta_{n}}^{\star}=\pi_{\alpha}^{\star}$ with $\pi^{\star}_{\alpha}\in\mathcal{P}_{2}^{ac}(\mathbb{R}^{d})$. For any $(\pi_{n})_{n\in\mathbb{N}^{\star}}\in\mathcal{P}_{2}^{ac}(\mathbb{R}^{d})^{\mathbb{N}^{\star}}$ such that $\lim_{n\to+\infty}\pi_{n}=\pi$ we have using Fatou’s lemma
@@
    This guarantees condition (a) in Definition 1 holds.
    To check condition (b) in Definition 1 consider w.l.o.g. $\pi\in\mathcal{P}_{2}^{ac}(\mathbb{R}^{d})$ which assigns positive mass to $1,1]^{d}$. For any $n\in\mathbb{N}^{\star}$, we denote $\pi_{n}\in\mathcal{P}_{2}^{ac}(\mathbb{R}^{d})$ with density w.r.t. the Lebesgue measure given for any $x\in\mathbb{R}^{d}$ by
@@
    Therefore, we have $\lim_{n\to+\infty}\mathbf{W}_{2}(\pi_{n},\pi)^{2}=0$ as $\pi_{n}\to\pi$ weakly and $\int_{\mathbb{R}^{d}}\left\|x\right\|^{2}\pi_{n}(x)\mathrm{d}x\to\int_{\mathbb{R}^{d}}\left\|x\right\|^{2}\pi(x)\mathrm{d}x$. We also observe that by construction $0<Z_{1}\leq Z_{2}\leq\dots\leq 1$ and $|\pi_{n}(x)|\leq\pi(x)/Z_{1}$.
    We now show that $\lim_{n\to+\infty}\mathcal{F}^{\eta_{n}}(\pi_{n})=\mathcal{F}(\pi)$. Consider the function $x\mapsto\pi_{n}(x)\log(\varphi(x)+\lambda\pi_{n}[\mathrm{k}(x,\cdot)]+\eta_{n})$. For any $n\in\mathbb{N}^{\star}$ and $x\in\mathbb{R}^{d}$, $\log(\varphi(x)+\lambda\pi_{n}[\mathrm{k}(x,\cdot)]+\eta_{n})\leq\log(2\lambda\mathtt{M}+\sup_{n\in\mathbb{N}}\eta_{n})$ and thus
    $\pi_{n}(x)\log(\varphi(x)+\lambda\pi_{n}[\mathrm{k}(x,\cdot)]+\eta_{n})\leq\log(2\lambda\mathtt{M}+\sup_{n\in\mathbb{N}}\eta_{n})\frac{\pi(x)}{Z_{1}}.$ (25)
    Define $\Phi:\ \mathbb{R}^{d}\to\mathbb{R}$, $\Phi:x\mapsto\left\|x\right\|^{2}$. Using Jensen’s inequality and the fact that under Assumption 2 there exists $C_{2}\geq 0$ such that for any $(x,y)\in\mathbb{R}^{d}\times\mathbb{R}^{d}$, $\exp[-\Phi(x)-\Phi(y)]/\mathrm{k}(x,y)\leq C_{2}\exp[C_{2}]$, we have for any $y\in\mathbb{R}^{d}$ and $n\in\mathbb{N}^{\star}$ that
@@
    $\geq\lambda\left(-\Phi(x)-\sup_{n\in\mathbb{N}^{\star}}\pi_{n}[\Phi]-C_{2}-\log C_{2}\right),$
    where we used the fact that $\varphi(x)\geq 0$ under Assumption 1. Since, $(\pi_{n})_{n\in\mathbb{N}^{\star}}$ is relatively compact in $\mathcal{P}_{2}(\mathbb{R}^{d})$ [4, 7.1.5], there exists $C\geq 0$ such that for any $n\in\mathbb{N}^{\star}$, $\pi_{n}[\Phi]\leq C$, see [56, Definition 6.8]. Hence,
    $\pi_{n}(x)\log(\varphi(x)+\lambda\pi_{n}[\mathrm{k}(x,\cdot)]+\eta_{n})\geq\lambda\left(-\Phi(x)-C-C_{2}-\log C_{2}\right)\frac{\pi(x)}{Z_{1}}.$ (26)
@@
    As $\pi\in\mathcal{P}_{2}^{ac}(\mathbb{R}^{d})$, its entropy $\mathrm{H}(\pi)$ is finite and from the definition of the Lebesgue integral we have that $\int_{\mathbb{R}^{d}}\pi(x)|\log\pi(x)|\mathrm{d}x<\infty$.
@@
    Therefore, we have that $\mathcal{F}_{\alpha}(\pi)=\lim_{n\to+\infty}\mathcal{F}_{\alpha}^{\eta_{n}}(\pi_{n})$, and $\{\mathcal{F}_{\alpha}^{\eta}:\,\eta\geq 0\}$ is a $\Gamma$-system. In addition, for any $\eta>0$, $\mathcal{F}_{\alpha}^{\eta}$ admits a unique minimizer $\pi_{\alpha,\eta}^{\star}$ by Proposition 1. Using Proposition 8, we get that $\pi_{\alpha}^{\star}$ is a minimizer of $\mathcal{F}_{\alpha}$.
@@
    By Lemma 2, $\pi_{\alpha_{n},\eta_{n}}^{\star}\in\mathcal{P}_{2}^{ac}(\mathbb{R}^{d})$. Consider $\{\mathcal{F}_{\alpha}^{\eta}:\alpha,\eta\geq 0\}$. Let $(\alpha_{n})_{n\in\mathbb{N}}\in(0,+\infty)^{\mathbb{N}},(\eta_{n})_{n\in\mathbb{N}}\in[0,+\infty)^{\mathbb{N}}$ be such that $\underset{n\to+\infty}{\lim}\alpha_{n}=0$, $\underset{n\to+\infty}{\lim}\eta_{n}=0$. For any $(\pi_{n})_{n\in\mathbb{N}}\in\mathcal{P}^{ac}(\mathbb{R}^{d})^{\mathbb{N}}$ such that $\lim_{n\to+\infty}\pi_{n}=\pi$ we have using the monotonicity in $\alpha$ of $\mathcal{F}_{\alpha}^{\eta}$ and Fatou’s lemma
@@
    This guarantees condition (a) in Definition 1 holds.
    To check condition (b) in Definition 1 consider for $\pi\in\mathcal{P}_{2}^{ac}(\mathbb{R}^{d})$ and for any $n\in\mathbb{N}$, $\pi_{n}\in\mathcal{P}_{2}^{ac}(\mathbb{R}^{d})$ defined as in (24). The proof of point 1 shows that $\mathcal{F}^{\eta_{n}}(\pi_{n})\to\mathcal{F}(\pi)$ as $n\to\infty$.
@@
    with $g(x)$ integrable since $\pi\in\mathcal{P}_{2}^{ac}(\mathbb{R}^{d})$ implies $\mathrm{H}(\pi)<\infty$.

### v3

@@ 2.1 Stability of $\mathcal{F}_{\alpha}$
Assumption 1 restricts the class of kernels $\mathrm{k}$ to smooth kernels with bounded derivatives up to the third order. This might seem restrictive, however, smooth kernels are particularly challenging while discountinuous or degenerate kernels normally allow for specific solution methodologies (see, e.g., [41, Ch. 11]). In this section, we will require only that $\mathrm{k}\in\mathrm{C}(\mathbb{R}^{d}\times\mathbb{R}^{d},[0,+\infty))$. The control of higher order derivatives is used to establish convergence result for the interacting particle system (in both continuous and discrete time) .
@@ 2.2 Convergence of Minimizers
We now consider the limiting behaviour of the regularized functional $\mathcal{F}_{\alpha}^{\eta}$ when the parameters $\alpha,\eta$ converge to 0. In particular, we show that if both parameters tend to 0 the minimizer of $\mathcal{F}_{\alpha}^{\eta}$ converges to one of the minimizers of the unregularized functional under some regularity conditions.
@@
Under Assumptions 1 and 2 we can show that any minimizer of $\mathcal{F}_{\alpha}^{\eta}$ for $\alpha>0,\ \eta\geq 0$ has finite second moment (Lemma 2), i.e. $\pi_{\alpha,\eta}^{\star}\in\mathcal{P}_{2}^{ac}(\mathbb{R}^{d})$.
Let $\mathcal{Q}_{2}^{ac}(\mathbb{R}^{d})$ denote the set of probability distributions with finite second moment, which admit an essentially bounded density w.r.t. the Lebesgue measure, i.e. $\pi\in\mathcal{P}_{2}^{ac}(\mathbb{R}^{d})$ and $\mathrm{ess}\sup_{x\in\mathbb{R}^{d}}\pi(x)<m$ for some $m>0$. Under the assumption that $\pi_{\alpha,\eta}^{\star}\in\mathcal{Q}_{2}^{ac}(\mathbb{R}^{d})$, we have the following results which describe the behaviour of the minimizer as $\eta\to 0$ and as $(\alpha,\eta)\to(0,0)$ jointly.
@@
Assume 1 and 2 and that $\pi_{\alpha,\eta}^{\star}\in\mathcal{Q}_{2}^{ac}(\mathbb{R}^{d})$ for all $\alpha>0,\ \eta\geq 0$.
@@
    If there exist $\pi_{\alpha}^{\star}\in\mathcal{Q}_{2}^{ac}(\mathbb{R}^{d})$, $(\eta_{n})_{n\in\mathbb{N}^{\star}}\in[0,+\infty)^{\mathbb{N}^{\star}}$ such that $\lim_{n\to+\infty}\eta_{n}=0$ and $\lim_{n\to+\infty}\mathbf{W}_{2}(\pi_{\alpha,\eta_{n}}^{\star},\pi_{\alpha}^{\star})=0$, then
    $\pi_{\alpha}^{\star}\in\arg\min_{\mathcal{Q}_{2}^{ac}(\mathbb{R}^{d})}\mathcal{F}_{\alpha}(\pi).$
@@
    If there exist $\pi^{\star}\in\mathcal{Q}_{2}^{ac}(\mathbb{R}^{d})$, $(\eta_{n})_{n\in\mathbb{N}^{\star}}\in[0,+\infty)^{\mathbb{N}^{\star}},(\alpha_{n})_{n\in\mathbb{N}^{\star}}\in(0,+\infty)^{\mathbb{N}}$ such that $\lim_{n\to+\infty}\eta_{n}=0,\lim_{n\to+\infty}\alpha_{n}=0$ and $\lim_{n\to+\infty}\mathbf{W}_{2}(\pi_{\alpha,\eta_{n}}^{\star},\pi^{\star})=0$, then
    $\pi^{\star}\in\arg\min_{\mathcal{Q}_{2}^{ac}(\mathbb{R}^{d})}\mathcal{F}(\pi).$
@@ 2.3 Maximum Entropy Methods
Maximum entropy solutions to Fredholm integral equations were originally proposed in [46] under moment constraints obtained by selecting different families of basis functions [37, 34]. In the one dimensional case, such a solution can be written analytically (see, e.g., [34, Prop. 3.1]), and these approaches are particularly effective. However, in higher dimensions maximum entropy methods generally require discretization of the support of the solution $\pi$.
@@ 3.2 McKean–Vlasov SDE
The previous result deals with the behaviour of $\pi_{t}$ for fixed $\alpha,\eta$. Under our assumptions, [14, Thm. 4.1] and [48, Thm. 1] guarantee that $\mathcal{F}^{\eta}(\pi_{t})-\inf\mathcal{F}^{\eta}\to 0$ provided that the parameter $\alpha$ decays at rate $1/\log t$.
@@ 5.1 Simple Gaussian Example
We start by considering a simple analytically-tractable integral equation. Consider, for $\lambda\in(0,1)$, (1) with $\varphi(x)=(1-\lambda)\mathcal{N}\left(x;0,1\right)$ and $\mathrm{k}(x,y)=\mathcal{N}\left(y;xe^{-\beta},1-e^{-2\beta}\right)$ for some $\beta>0$, with $\mathcal{N}(x;\mu,\sigma^{2})$ denoting the density of a normal distribution of mean $\mu$ and variance $\sigma^{2}$ evaluated at $x$. Since $\max_{x}\int_{\mathbb{R}^{d}}\lambda k(x,y)\mathrm{d}y=\lambda<1$, [41, Cor. 2.16] guarantees that (1) admits a unique solution, $\pi(x)=\mathcal{N}\left(x;0,1\right)$ for all $0<\lambda<1$. In our experiments we set $\beta=0.5$.
@@ 5.1.2 Comparison with Reversible jump MCMC
The previous experiments show that the presence of the reference measure significantly improves the results provided by FE2kind-WGF. We now compare our approach with the MCMC algorithm of [23] which provides a Monte Carlo approximation of the von Neumann series representation of the solution $\pi$ (see, e.g., [41, Sec. 2.4]). We implement their reversible jump MCMC (RJ-MCMC) with the following set up: we set the death and birth probabilities to 1/3 and use a random walk proposal with variance $\sigma^{2}=0.1^{2}$ for the update move and sample from $\pi$ for the birth move.
@@ 5.2 Karhunen–Loève Expansions
The Karhunen–Loève expansion of a given stochastic process with covariance function $\mathrm{k}(x,y)$ is given by the pairs of eigenvalues and eigenfunctions $(\lambda,\pi)$ of $\mathrm{k}$. For any eigenvalue $\lambda$, the corresponding eigenfunction can be found by solving (1) with $\varphi\equiv 0$. We consider two examples commonly used in the Gaussian process literature [61, Sec. 4.2]: the squared exponential kernel, $\mathrm{k}(x,y)=\exp(-(y-x)^{2})$, which satisfies our assumptions, and the exponential kernel, $\mathrm{k}(x,y)=\exp(-|y-x|)$, which is not differentiable for $y=x$ and therefore does not satisfy Assumption 1.
@@ A.2 Proof of Proposition 1
    $\mathcal{F}^{\eta}(\pi)\geq$ $-e^{-1}+\frac{\eta}{(1+\lambda)\mathtt{M}+\eta}\mathcal{F}^{\eta}(\pi)$
    and thus
    $\mathcal{F}^{\eta}(\pi)\geq-\left(1-\frac{\eta}{(1+\lambda)\mathtt{M}+\eta}\right)^{-1}e^{-1}=:\mathtt{C}.$
@@
    $\mathcal{F}_{\alpha}^{\eta}(\pi_{0})=-\mathrm{H}(\pi_{0})-\int_{\mathbb{R}^{d}}\pi_{0}(x)\log(\varphi(x)+\lambda\pi_{0}[\mathrm{k}(x,\cdot)]+\eta)\mathrm{d}x.$
@@
    $|\log(\varphi(x)+\lambda\pi[\mathrm{k}(x,\cdot)]+\eta)|\leq\max(|\log((1+\lambda)\mathtt{M}+\eta)|,|\log(\eta)|),$
@@
    $-\int_{\mathbb{R}^{d}}\pi_{0}(x)\log(\varphi(x)+\lambda\pi_{0}[\mathrm{k}(x,\cdot)]+\eta)\mathrm{d}x\leq\max(|\log((1+\lambda)\mathtt{M}+\eta)|,|\log(\eta)|).$
@@ B.2 Proof of Proposition 2
Since $t\mapsto t|\log(t)|$ is bounded on $[0,1]$ we have that
@@
    Consider for fixed $\alpha$ the family $\{\mathcal{F}_{\alpha}^{\eta}:\mathcal{Q}_{2}^{ac}(\mathbb{R}^{d})\to\mathbb{R}^{+},\ s.t.\ \eta\geq 0\}$. Let $(\eta_{n})_{n\in\mathbb{N}^{\star}}\in[0,+\infty)^{\mathbb{N}^{\star}}$ be such that $\underset{n\to+\infty}{\lim}\eta_{n}=0$. For any $(\pi_{n})_{n\in\mathbb{N}^{\star}}\in\mathcal{P}_{2}^{ac}(\mathbb{R}^{d})^{\mathbb{N}^{\star}}$ such that $\lim_{n\to+\infty}\pi_{n}=\pi$ we now show that
@@
    Using the properties of $\liminf$ and the lower semi-continuity of the KL divergence [24, Lemma 1.4.3-(b)] we have
    $\liminf_{n\to+\infty}\mathcal{F}_{\alpha}^{\eta_{n}}(\pi_{n})$ $\geq\liminf_{n\to+\infty}\mathcal{F}^{\eta_{n}}(\pi_{n})+\alpha\liminf_{n\to+\infty}\textup{KL}\left(\pi_{n}\middle|\pi_{0}\right)$
    $\geq\liminf_{n\to+\infty}\mathcal{F}^{\eta_{n}}(\pi_{n})+\alpha\textup{KL}\left(\pi\middle|\pi_{0}\right).$
    Let us denote $\bar{\eta}:=\sup_{n\in\mathbb{N}}\eta_{n}$, then
    $\mathcal{F}^{\eta_{n}}(\pi_{n})$ $=\mathcal{F}^{\bar{\eta}}(\pi_{n})+\int_{\mathbb{R}^{d}}\pi_{n}(x)\log\frac{\varphi(x)+\lambda\pi_{n}[\mathrm{k}(x,\cdot)]+\bar{\eta}}{\varphi(x)+\lambda\pi_{n}[\mathrm{k}(x,\cdot)]+\eta_{n}}\mathrm{d}x.$
    Using again the properties of $\liminf$ and the lower semi-continuity of $\mathcal{F}^{\eta}$ established in Proposition 1 we have
    $\liminf_{n\to+\infty}\mathcal{F}^{\eta_{n}}(\pi_{n})$ $\geq\mathcal{F}^{\bar{\eta}}(\pi)+\liminf_{n\to+\infty}\int_{\mathbb{R}^{d}}\pi_{n}(x)\log\frac{\varphi(x)+\lambda\pi_{n}[\mathrm{k}(x,\cdot)]+\bar{\eta}}{\varphi(x)+\lambda\pi_{n}[\mathrm{k}(x,\cdot)]+\eta_{n}}\mathrm{d}x.$
    Moreover, the function
    $f_{n}(x):=\pi_{n}(x)\log\frac{\varphi(x)+\lambda\pi_{n}[\mathrm{k}(x,\cdot)]+\bar{\eta}}{\varphi(x)+\lambda\pi_{n}[\mathrm{k}(x,\cdot)]+\eta_{n}}\geq 0$
    by construction and therefore Fatou’s Lemma allows us to conclude
    $\liminf_{n\to+\infty}\mathcal{F}^{\eta_{n}}(\pi_{n})$ $\geq\mathcal{F}^{\bar{\eta}}(\pi)$
    $+\int_{\mathbb{R}^{d}}\pi(x)\log\frac{\varphi(x)+\lambda\pi_{n}[\mathrm{k}(x,\cdot)]+\bar{\eta}}{\varphi(x)+\lambda\pi_{n}[\mathrm{k}(x,\cdot)]}\mathrm{d}x$
    $=\mathcal{F}(\pi).$
    This guarantees condition (a) in Definition 1 holds for any $(\pi_{n})_{n\in\mathbb{N}^{\star}}\in\mathcal{Q}_{2}^{ac}(\mathbb{R}^{d})^{\mathbb{N}^{\star}}\subset\mathcal{P}_{2}^{ac}(\mathbb{R}^{d})^{\mathbb{N}^{\star}}$.
    To check condition (b) in Definition 1 consider w.l.o.g. $\pi\in\mathcal{Q}_{2}^{ac}(\mathbb{R}^{d})$ which assigns positive mass to $1,1]^{d}$. For any $n\in\mathbb{N}^{\star}$, we consider $\pi_{n}$ with density w.r.t. the Lebesgue measure given for any $x\in\mathbb{R}^{d}$ by
@@
    We observe that by construction $0<Z_{1}\leq Z_{2}\leq\dots\leq 1$ and $|\pi_{n}(x)|\leq\pi(x)/Z_{1}$ and so $\pi_{n}\in\mathcal{Q}_{2}^{ac}(\mathbb{R}^{d})$ for all $n\geq 1$. Therefore, we have $\lim_{n\to+\infty}\mathbf{W}_{2}(\pi_{n},\pi)^{2}=0$ as $\pi_{n}\to\pi$ weakly and $\int_{\mathbb{R}^{d}}\left\|x\right\|^{2}\pi_{n}(x)\mathrm{d}x\to\int_{\mathbb{R}^{d}}\left\|x\right\|^{2}\pi(x)\mathrm{d}x$.
    We now show that $\lim_{n\to+\infty}\mathcal{F}^{\eta_{n}}(\pi_{n})=\mathcal{F}(\pi)$. Consider the function $x\mapsto\pi_{n}(x)\log(\varphi(x)+\lambda\pi_{n}[\mathrm{k}(x,\cdot)]+\eta_{n})$. For any $n\in\mathbb{N}^{\star}$ and $x\in\mathbb{R}^{d}$, $\log(\varphi(x)+\lambda\pi_{n}[\mathrm{k}(x,\cdot)]+\eta_{n})\leq\log((1+\lambda)\mathtt{M}+\sup_{n\in\mathbb{N}}\eta_{n})$ and thus
    $\pi_{n}(x)\log(\varphi(x)+\lambda\pi_{n}[\mathrm{k}(x,\cdot)]+\eta_{n})\leq\log((1+\lambda)\mathtt{M}+\sup_{n\in\mathbb{N}}\eta_{n})\frac{\pi(x)}{Z_{1}}.$ (25)
    Define $\Phi:\ \mathbb{R}^{d}\to\mathbb{R}$, $\Phi:x\mapsto\left\|x\right\|^{2}$. Using Jensen’s inequality and the fact that under Assumption 2 there exists $C_{2}\geq 0$ such that for any $(x,y)\in\mathbb{R}^{d}\times\mathbb{R}^{d}$, $\mathrm{k}(x,y)\geq C_{2}^{-1}\exp[-C_{2}(1+\Phi(x)+\Phi(y))]$, we have for any $y\in\mathbb{R}^{d}$ and $n\in\mathbb{N}^{\star}$ that
@@
    $\geq\lambda\left(-C_{2}\Phi(x)-C_{2}\sup_{n\in\mathbb{N}^{\star}}\pi_{n}[\Phi]-C_{2}-\log C_{2}\right),$
    where we used the fact that $\varphi(x)\geq 0$ under Assumption 1. Since, $(\pi_{n})_{n\in\mathbb{N}^{\star}}$ is relatively compact in $\mathcal{P}_{2}(\mathbb{R}^{d})$ [4, Proposition 7.1.5], there exists $C\geq 0$ such that for any $n\in\mathbb{N}^{\star}$, $\pi_{n}[\Phi]\leq C$, see [59, Definition 6.8]. Hence,
    $\pi_{n}(x)\log(\varphi(x)+\lambda\pi_{n}[\mathrm{k}(x,\cdot)]+\eta_{n})\geq\lambda\left(-C_{2}\Phi(x)-C_{2}(C+1)-\log C_{2}\right)\frac{\pi(x)}{Z_{1}}.$ (26)
@@
    As $\pi\in\mathcal{Q}_{2}^{ac}(\mathbb{R}^{d})\subset\mathcal{P}_{2}^{ac}(\mathbb{R}^{d})$ it has finite second moment. This and the fact that $\pi$ has essentially bounded density imply that its entropy $\mathrm{H}(\pi)$ is finite by [30, Theorem 1]. From the definition of the Lebesgue integral we have that $\int_{\mathbb{R}^{d}}\pi(x)|\log\pi(x)|\mathrm{d}x<\infty$.
@@
    Therefore, we have that $\mathcal{F}_{\alpha}(\pi)=\lim_{n\to+\infty}\mathcal{F}_{\alpha}^{\eta_{n}}(\pi_{n})$, and $\{\mathcal{F}_{\alpha}^{\eta}:\mathcal{Q}_{2}^{ac}(\mathbb{R}^{d})\to\mathbb{R}^{+},\ s.t.\,\eta\geq 0\}$ is a $\Gamma$-system. In addition, for any $\eta>0$, $\mathcal{F}_{\alpha}^{\eta}$ admits a unique minimizer $\pi_{\alpha,\eta}^{\star}$ by Proposition 1, which under our assumptions, satisfies $\pi_{\alpha,\eta}^{\star}\in\mathcal{Q}_{2}^{ac}(\mathbb{R}^{d})$. Using Proposition 8, we get that $\pi_{\alpha}^{\star}$ is a minimizer of $\mathcal{F}_{\alpha}$.
@@
    Consider $\{\mathcal{F}_{\alpha}^{\eta}:\mathcal{Q}_{2}^{ac}(\mathbb{R}^{d})\to\mathbb{R}^{+},\ s.t.\ \alpha,\eta\geq 0\}$. Let $(\alpha_{n})_{n\in\mathbb{N}}\in(0,+\infty)^{\mathbb{N}},(\eta_{n})_{n\in\mathbb{N}}\in[0,+\infty)^{\mathbb{N}}$ be such that $\underset{n\to+\infty}{\lim}\alpha_{n}=0$, $\underset{n\to+\infty}{\lim}\eta_{n}=0$. For any $(\pi_{n})_{n\in\mathbb{N}}\in\mathcal{P}^{ac}(\mathbb{R}^{d})^{\mathbb{N}}$ such that $\lim_{n\to+\infty}\pi_{n}=\pi$, using the monotonicity in $\alpha$ of $\mathcal{F}_{\alpha}^{\eta}$ and the result in point 1. we have
@@
    This guarantees condition (a) in Definition 1 holds since $\mathcal{Q}_{2}^{ac}(\mathbb{R}^{d})\subset\mathcal{P}^{ac}(\mathbb{R}^{d})$.
    To check condition (b) in Definition 1 consider for $\pi\in\mathcal{Q}_{2}^{ac}(\mathbb{R}^{d})$ and for any $n\in\mathbb{N}$, $\pi_{n}\in\mathcal{Q}_{2}^{ac}(\mathbb{R}^{d})$ defined as in (24). The proof of point 1 shows that $\mathcal{F}^{\eta_{n}}(\pi_{n})\to\mathcal{F}(\pi)$ as $n\to\infty$.
@@
    with $g(x)$ integrable since $\pi\in\mathcal{Q}_{2}^{ac}(\mathbb{R}^{d})$ implies $\mathrm{H}(\pi)<\infty$ by [30, Theorem 1].